# Isomorphism classes of generating sets

### v1

@@ 3. Finite support D-based iteration
If $e<_{D}d$ then
$\dot{U}_{<e}\subseteq\dot{U}_{e}\subseteq\dot{U}_{<d}\subseteq\dot{U}_{d}$
is a tower of ultrafilters of the models
$V^{\mathbb{P}_{<e}}\subseteq V^{\mathbb{P}_{e}}\subseteq V^{\mathbb{P}_{<d}}\subseteq V^{\mathbb{P}_{d}}$
Indeed, $\dot{U}_{<e}\subseteq\dot{U}_{e}\subseteq\dot{U}_{<d}$. Since the generic set $\dot{B}_{d}$ added by ˙$\dot{\mathbb{Q}}_{d}$ diagonalizes $\dot{U}_{<d}$, and $\dot{B}_{d}\in\dot{U}_{d}$, it follows that $\dot{U}_{<d}\subseteq\dot{U}_{d}$.
@@
Assume that $\mathbb{D}$ is well-founded. Then the set of conditions $p\in\mathbb{M}_{\mathbb{D}}$ such that for all $d\in\operatorname{Supp}(p)$, $p_{d}=(\check{\tau}_{d},\dot{A}_{d})$ is dense.
@@
This is an easy induction which uses the fact that $\mathbb{M}_{\mathbb{D}}$ has finite support. ∎
Lemma 3.6.
For every regular uncountable cardinal $\lambda$, $\mathbb{M}_{\mathbb{D}}$ is $\lambda$-Knaster.
Proof.
This is a straightforward application of the previous lemma combined with the $\Delta$-system lemma and the fact that Mathias forcing is $\sigma$-centered. ∎
Let $G\subseteq\mathbb{M}_{\mathbb{D}}$ be a generic. For every $d\in D$, denote $B_{d}=(\dot{B}_{d})_{G}$.
@@
($(\{B_{d}\mid d\in D\},\supseteq^{*})$ is isomorphic to $D$.
Proof.
Let us begin observing that if $e<_{D}d$ then $B_{d}\subseteq^{*}B_{e}$. To see this, note that $B_{e}\in U_{e}\subseteq U_{<d}$ by Remark 3.4. Moreover $B_{d}$ diagonalizes $U_{<d}$, and so $B_{d}\subseteq^{*}B_{e}$.
Now we will show that if $e\nleq_{D}d$, then $B_{d}\not\subseteq^{*}B_{e}$. Let $D^{\prime}=D_{\leq d}\cup D_{<e}$. Since $D^{\prime}$ is downwards closed, $G$ projects onto a $\mathbb{M}_{\mathbb{D}^{\prime}}$-generic $G^{\prime}$ by Lemma 3.1. Towards a contradiction, suppose there is a condition $p\in\mathbb{P}_{D^{\prime}\cup\{e\}}$ forcing ``$\dot{B}_{d}\setminus\check{m}\subseteq\dot{B}_{e}"$. (Notice that $\dot{B}_{d}$ is a $\mathbb{P}_{D^{\prime}}$-name). Extend $p\restriction D^{\prime}$ to a condition $p^{\prime}\in\mathbb{P}_{D^{\prime}}$ which forces for some $k<\omega$ that ``$\check{k}\in\dot{B}_{d}\setminus\max(\check{s}\cup\{\check{m}\})"$, where $p(e)=(\check{s},\dot{A})$. Let $p^{*}\in\mathbb{P}_{D^{\prime}\cup\{e\}}$ be the condition defined by $p^{*}\restriction D^{\prime}=p^{\prime}$ and $p^{*}(e)=\langle\check{s},\dot{A}\setminus(\check{k}+1)\rangle$ and notice that $p^{*}\leq_{D^{\prime}\cup\{e\}}p$. However, $p^{*}$ forces the negation of ``$\check{B}_{d}\setminus\check{m}\subseteq\dot{B}_{e}"$, which is impossible.
Corollary 3.8.
[…]
Proof.
Since each $B_{d}$ diagonalizes $U_{<d}$, the “moreover” part follows from the equality $\bigcup_{d\in D}U_{d}=\bigcup_{d\in D}U_{<d}$, which follows from Remark 3.4 and the fact that $D$ has no maximal elements.
To see the first part, note that $U$ has the finite intersection property since it is the directed union of a family of sets with the finite intersection property, namely the $U_{d}$’s.
To show that $U$ is an ultrafilter, it suffices to show that for every $X\in P(\omega)^{V[G]}$, either $X\in U$ or $\omega\setminus X\in U$. Since each condition in $\mathbb{M}_{\mathbb{D}}$ has finite support, and since $\mathbb{M}_{\mathbb{D}}$ has the ccc, whenever $X\subseteq\omega$, there is a nice name $\dot{X}$ based on maximal antichains $A_{n}$ for $n<\omega$. Let $\sigma=\bigcup\{\operatorname{Supp}(p)\mid p\in A_{n},\ n<\omega\}\subseteq D$. By our assumption, $D$ is countably directed, let $d\in D$ be an upper bound for $\sigma$. Thus $X\in V^{\mathbb{P}_{\leq d}}$ and since $U_{d}$ is an ultrafilter of $V^{\mathbb{P}_{\leq d}}$, either $X\in U_{d}$ or $\omega\setminus X\in U_{d}$. Since $U_{d}\subseteq U$, either $X\in U$ or $\omega\setminus X\in U$. ∎
Finally, we would like to construct, for a restricted choice of $\mathbb{D}$, a suitable sequence of names $\langle\dot{U}_{<d},\dot{U}_{d}:d\in D\rangle$.
Theorem 3.9.
[…]
Proof.
We use the well foundedness to define $\dot{U}_{<d}$ and $\dot{U}_{d}$ recursively. First, we let $\dot{U}_{d}$ be a canonical name for any uniform ground model ultrafilter on $\omega$ whenever $d$ is $D$-minimal.
By II, once $\dot{U}_{<d}$ was defined, $\dot{B}_{d}$ is forced to diagonalize $\dot{U}_{<d}$. Hence, the set $\dot{U}_{<d}\cup\{\dot{B}_{d}\}$ is forced to have the finite intersection property and therefore can be extended (arbitrarily) to an ultrafilter. Let $\dot{U}_{d}$ be a name for such an ultrafilter. The non-trivial part of the construction is to guarantee condition I, suppose that $\langle\dot{U}_{<e},\dot{U}_{e}:e<_{D}d\rangle$ has been defined.
Lemma 3.10.
$\Vdash_{\mathbb{P}_{<d}}``\bigcup_{e<_{D}d}\dot{U}_{e}$ has the finite intersection property”.
Proof.
We prove the lemma by induction on the number of sets intersected. Specifically, for each $k\geq 1$ we show that if $e_{0},\dots,e_{k}<_{D}d$, and $Y_{i}\in U_{e_{i}}$, then $\bigcap_{i=0}^{k}Y_{i}$ is infinite. For $k=1$, let $Y_{0}\in U_{e_{0}}$ and $Y_{1}\in U_{e_{1}}$. Towards contradiction, suppose that $Y_{0}\cap Y_{1}$ is finite and let $p\in\mathbb{P}_{<d}$ force that $\dot{Y}_{0}\cap\dot{Y}_{1}\setminus l=\emptyset$ for some $l<\omega$. Let $e^{m}=e_{0}\wedge e_{1}$ be the meet of $e_{0},e_{1}$. Let $G_{\leq e^{m}}$ be a $V$-generic filter for $\mathbb{P}_{\leq e^{m}}$. Working in $V[G_{\leq e^{m}}]$, define:
$\tilde{A}=\Big{\{}n<\omega\mid\exists q\in\mathbb{P}_{\leq e_{0}}/G_{\leq e^{m}},\ q\leq p\restriction D_{\leq e_{0}},\ q\Vdash_{\mathbb{P}_{\leq e_{0}}}\check{n}\in\dot{Y}_{0}\Big{\}}$
$\tilde{B}=\Big{\{}n<\omega\mid\exists q\in\mathbb{P}_{\leq e_{1}}/G_{\leq e^{m}},\ q\leq p\restriction D_{\leq e_{1}},\ q\Vdash_{\mathbb{P}_{\leq e_{1}}}\check{n}\in\dot{Y}_{1}\Big{\}}$
We claim that $\tilde{A}\cap\tilde{B}\setminus l=\emptyset$. Otherwise, let $n\in\tilde{A}\cap\tilde{B}\setminus l$ and let $q_{0},q_{1}$ witness this. Since the conditions $q_{0},q_{1}$ are in quotients over the same forcing, there is $r\in\mathbb{P}_{\leq e^{m}}$ such that
$r\leq q_{0}\restriction D_{\leq e^{m}},q_{1}\restriction D_{\leq e^{m}}\leq p\restriction D_{\leq e^{m}}.$
Define $p^{*}\in\mathbb{P}_{<d}$ as follows, for $x<_{D}d$, set
$p^{*}(x)=\begin{cases}r(x)&x\leq_{D}e^{m}\\ q_{0}(x)&x\in D_{\leq e_{0}}\setminus D_{\leq e^{m}}\\ q_{1}(x)&x\in D_{\leq e_{1}}\setminus D_{\leq e^{m}}\\ p(x)&\text{otherwise}\end{cases}.$
Note that there is no conflict in the definition since if $x\leq_{D}e_{0}$ and $x\leq_{D}e_{1}$ then $x\leq_{D}e_{0}\wedge e_{1}=e^{m}$. We conclude that $p^{*}\leq p$ and $p^{*}\Vdash_{\mathbb{P}_{<d}}n\in\dot{Y}_{0}\cap\dot{Y}_{1}\setminus l$, contradiction.
However, by Remark 3.4, $\tilde{A}\in U_{e_{0}}\cap V[G_{\leq e^{m}}]=U_{e^{m}}$ and $\tilde{B}\in U_{e_{1}}\cap V[G_{\leq e^{m}}]=U_{e^{m}}$. Thus, $\tilde{A},\tilde{B}\in U_{e^{m}}$, producing a contradiction to the choice of $U_{e^{m}}$ being a uniform ultrafilter.
Now for the general case, suppose that $k\geq 2$. If there are $i\neq j$ such that $e_{i}$ and $e_{j}$ are $<_{D}$-comparable, then we may reduce the $k+1$ sets $Y_{0},..,Y_{k}$ to $k$ sets and apply the induction hypothesis for $k-1$. Suppose that $e_{0},...,e_{k}$ are pairwise incomparable. Since $D$ has degree $2$, there are $i\neq j$ and $e^{M}<_{D}d$ such that $e_{i},e_{j}\leq_{D}e^{M}$. Also, since $e_{i}$ and $e_{j}$ are incomparable, $e_{i},e_{j}<_{D}e^{M}$. Therefore, $U_{e_{i}},U_{e_{j}}\subseteq U_{<e^{M}}$. It follows that $B_{e^{M}}\in U_{e^{M}}$ is $\subseteq^{*}$-below both $Y_{i}$ and $Y_{j}$. Replacing $Y_{i},Y_{j}$ by $B_{e^{M}}$, we may now apply the induction hypothesis for $k-1$ elements. This concludes the inductive proof.
Corollary 3.11.
Suppose that $\mathbb{D}$ is a well-founded, degree $2$, countably directed semi-lattice, with no maximal elements, then there is a ccc forcing extension in which there is a uniform simple $P_{\mathbb{D}}$-point ultrafilter $U$.
Corollary 3.12.
Let $\lambda_{1},\lambda_{2}$ be any cardinals of uncountable cofinality. There is a ccc forcing extension in which there is a uniform $P_{\lambda_{1}\times\lambda_{2}}$-point ultrafilter.
Corollary 3.13.
Let $\mathbb{D}$ be a well-founded countably-directed two-dimensional order with no maximal elements. Then there is a ccc forcing extension in which there is a uniform $P_{\mathbb{D}}$-point ultrafilter.
Proof.
By Corollary 2.10, $\mathbb{D}$ can be cofinally embedded into a degree-two subset $A$ of a product of cardinals. Also since $\mathbb{D}$ is countably directed and has no maximal elements, so does $A$. Now we can apply Corollary 3.11 to force an ultrafilter $U$ which is a simple $P_{A}$-point. Since $\mathbb{D}$ cofinally embeds into $A$, by Proposition 2.1 $U$ is a simple $P_{\mathbb{D}}$-point. ∎
4. At a measurable cardinal
In this section, we produce examples of measures, i.e. normal ultrafilters, on a measurable cardinal $\kappa$ which have complicated yet controlled cofinal structures. As in Section 3 the idea is to produce a measure $U$ such that some poset with a complicated Tukey structure is embedded cofinally into $(U,\supseteq^{*})$, and we produce $U$ by a nonlinear iteration of a version of ultrafilter Mathias forcing. However, we will use a different method to choose the forcing posets used at each stage, which permits us to embed a more general class of posets into $(U,\supseteq^{*})$. For definiteness we will show how to embed the product of a finite increasing sequence of regular cardinals greater than $\kappa$.
@@ 4. At a measurable cardinal
To illustrate the main idea in a simple setting, suppose that $\kappa$ is a Laver indestructible supercompact cardinal and consider a linear iteration of ultrafilter Mathias forcing at $\kappa$ with $<\kappa$-supports, where at stage $\alpha$ we use a measure $U_{\alpha}\in V[G_{\alpha}]$ to define the next step in the iteration: there is no problem in choosing some $U_{\alpha}$ because the forcing so far is $\kappa$-directed closed. In the final model $V[G]$ there is of course a measure $U$, but a priori we have no reason to believe that the set $B_{\alpha}$ added at stage $\alpha$ is in $U$, or even that $U_{\alpha}\subseteq U$.

### v2

@@ 4. Mathias iteration
  1. (I)
    Every canonical $\mathbb{P}$-name for a subset of $\kappa$ is a canonical ${\mathbb{P}}_{a}$-name for some $a\in\mathbb{D}^{+}$. It follows that if $G$ is $\mathbb{P}$-generic then
    $P(\kappa)\cap V[G]=\bigcup_{a\in\mathbb{D}^{+}}P(\kappa)\cap V[G_{<a}]$
    where $G_{<a}$ is the projection of $G$ to ${\mathbb{P}}_{<a}$.
  2. (II)
    Similarly,
    1. (IIa)
      if $\beta<\mu^{+}$ and ${\rm cf}(\beta)>\kappa$, then
      $P(\kappa)\cap V[G_{\beta}]=\bigcup_{\alpha<\beta}P(\kappa)\cap V[G_{\alpha}].$
    2. (IIb)
      for any $\beta<\mu^{+}$,
      $P(\kappa)\cap V[G_{\beta}]=\bigcup_{d\in\mathbb{D}}P(\kappa)\cap V[G_{<(\beta,d)}]$
      In particular, if $\beta$ is an active stage, then
      $A_{\beta}[G_{\beta}]=\bigcup_{d\in\mathbb{D}}\dot{F}_{(\beta,d)}[G_{<(\beta,d)}].$
Let $\dot{U}$ be a canonical $\mathbb{P}$-name for a $\kappa$-complete ultrafilter over $\kappa$, and for $\beta<\mu^{+}$ consider the following $\mathbb{P}_{\beta\times\mathbb{D}}$-name
$\dot{U}_{\beta}=\{(q,\dot{\tau})\in\dot{U}:$ $q\in{\mathbb{P}}_{\beta\times\mathbb{D}}\text{ and}$
[…]
Then:
  • $\dot{U}\subseteq H_{\mu^{+}}$.
  • $\dot{U}=\bigcup_{\beta<\mu^{+}}{\dot{U}}_{\beta}$, and $\langle\dot{U}_{\beta}\mid\beta<\mu^{+}\rangle$ is continuous at points of cofinality greater than $\kappa$.
  • For all but non-stationary many $\beta\in\mu^{+}\cap\operatorname{cof}(\mu)$, $\dot{U}_{\beta}=\dot{U}\cap M_{\beta}$ is a name for a $\kappa$-complete ultraiflter on $\kappa$, and is forced to equal $\dot{U}\cap V[\dot{G}_{\beta\times\mathbb{D}}]$. Denote the set of such $\beta$ by $S_{\dot{U}}$.
Consider the following crucial set
$S^{*}_{\dot{U}}:=\{\beta\in S_{\dot{U}}:A_{\beta}=\dot{U}_{\beta}\},$
then by the diamond assumption, $S^{*}_{\dot{U}}$ is a stationary subset of $\mu^{+}\cap\operatorname{cof}(\mu)$, and clearly every $\beta\in S^{*}_{\dot{U}}$ is an “active stage” in the definition of $\mathbb{P}$. Note that for every $\beta\in S^{*}_{\dot{U}}$:
($\star$) $\dot{U}[G]\cap V[G_{\beta}]=\dot{U}_{\beta}[G_{\beta}]=A_{\beta}[G_{\beta}]\in V[G_{\beta}].$
Claim 4.3.
Let $\dot{U}$ be a canonical $\mathbb{P}$-name for a $\kappa$-complete ultrafilter over $\kappa$, let $\alpha\in S^{*}_{\dot{U}}$, and $G_{\alpha}\subseteq\mathbb{P}_{\alpha\times\mathbb{D}}$ be $V$-generic. Then:
  1. (1)
    For all $\beta\in S^{*}_{\dot{U}}\cap\alpha$, $\dot{U}_{\alpha}[G_{\alpha}]\cap V[G_{\beta}]=\dot{U}_{\beta}[G_{\beta}]$
  2. (2)
    $\dot{U}_{\alpha}[G_{\alpha}]=\bigcup_{d\in\mathbb{D}}\dot{F}_{(\alpha,d)}[G_{<(\alpha,d)}]$.
Proof.
$(1)$ follows straightforwardly from ($\star$4). For $(2)$, by the definition of $\dot{F}_{\alpha,d}$, and since $\dot{U}_{\alpha}=A_{\alpha}$, the inclusion from right to left is immediate. Let $x={\dot{x}}[G_{\alpha}]\in\dot{U}_{\alpha}[G_{\alpha}]$, let $p\in G_{\alpha}$ force $\dot{x}\in{\dot{U}}_{\alpha}$. Using the maximum principle and the fact that any name is forced to be equal to a canonical name, we find a canonical ${\mathbb{P}}_{\alpha\times\mathbb{D}}$ name $\dot{y}$ such that for any generic $H$, ${\dot{y}}[H]={\dot{x}}[H]$ when $p\in H$ and ${\dot{y}}[H]=\kappa$ when $p\notin H$. Then $x={\dot{y}}[G_{\alpha}]$ and $\Vdash_{\mathbb{P}_{\alpha\times\mathbb{D}}}\dot{y}\in\dot{U}_{\alpha}$. By the $\kappa^{+}$-cc and the $\kappa^{+}$-directedness of $\mathbb{D}$, we can find $d\in\mathbb{D}$ such that $\dot{y}$ is a $\mathbb{P}_{<(\alpha,d)}$-name. It follows that $x\in\dot{F}_{(\alpha,d)}[G_{<(\alpha,d)}]$. ∎
The following lemma gives a sufficient condition for the main theorem:
@@
Let $\dot{U}$ be a canonical $\mathbb{P}$-name for a $\kappa$-complete ultrafilter over $\kappa$ and let $I$ be a subset of $\mu^{+}$ such that $\{B_{(\gamma,d)}\mid\gamma\in I,d\in\mathbb{D}\}$ generates $\dot{U}[G]$ for some $V$-generic $G\subseteq\mathbb{P}$. Then in $V$ there is a club $C\subseteq\mu^{+}$, such that for every $\beta\in S^{*}_{\dot{U}}\cap C$:
  1. (1)
    $\{B_{(\gamma,d)}\mid\gamma\in I\cap\beta,d\in\mathbb{D}\}$ generates $\dot{U}_{\beta}[G_{\beta}]$.
  2. (2)
    If $I\times\mathbb{D}$ satisfies ($\dagger$3.5) of Corollary 3.5, and there is an unbounded $I^{*}\subseteq I\cap\beta$ with $\text{otp}(I^{*})=\mu$, then the ultrafilter $\dot{U}_{\beta}[G_{\beta}]$ is a simple $\mathbb{P}_{\mathbb{D}}$-point.
@@
For each $\alpha<\mu^{+}$, let $\beta_{\alpha}<\mu^{+}$ be the least such that for every $x\in\dot{U}[G]\cap V[G_{\alpha}]$ there is $\gamma\in\beta_{\alpha}\cap I$ and $d\in\mathbb{D}$ such that $B_{(\gamma,d)}\subseteq^{*}x$. Such a $\beta_{\alpha}$ exists since in $V[G_{\alpha}]$, $2^{\kappa}\leq\mu$. Let $C\subseteq\mu^{+}$ be the club of closure points of $\alpha\mapsto\beta_{\alpha}$. By the $\kappa^{+}$-cc of the forcing, we may assume that $C\in V$. To see $(1)$, let $\beta\in S^{*}_{\dot{U}}\cap C$, $x\in\dot{U}_{\beta}[G_{\beta}]$. By Remark 4.2(II), there is $\alpha\in S^{*}_{\dot{U}}\cap\beta$ such that $x\in\dot{U}_{\alpha}[G_{\alpha}]$. Since $\beta\in C$, $\beta_{\alpha}<\beta$ and therefore there is $\gamma\in I\cap\beta$ and $d\in\mathbb{D}$ such that $B_{(\gamma,d)}\subseteq^{*}x$.
For $(2)$, recall that $\langle d_{\alpha}\mid\alpha<\mu\rangle$ is a linearization of $\mathbb{D}$ and let $\langle\epsilon_{\alpha}\mid\alpha<\mu\rangle$ be the increasing enumeration of the members of $I^{*}$. By Corollary 3.5, for every $\alpha,\alpha^{\prime}<\mu$, and $a,b\in\mathbb{D}$, $B_{\epsilon_{\alpha},a}\subset^{*}B_{\epsilon_{\alpha^{\prime}},b}$ if and only if ${\alpha^{\prime}}<\alpha$ and $b<_{\mathbb{D}}a$. Let
$\mathcal{B}=\{B_{(\epsilon_{\alpha},d_{\alpha})}\mid\alpha<\mu\}.$
Then $(\mathcal{B},\supset^{*})\simeq(\{(\epsilon_{\alpha},d_{\alpha})\mid\alpha<\mu\},<_{\mathbb{D}^{+}})\simeq\mathbb{D}$, where the right isomorphism is given by $\varphi(\epsilon,d)=d$. While this is clearly a bijection, let us varify that is is order preserving: let $\alpha,{\alpha^{\prime}}<\mu$. Then
$(\epsilon_{\alpha^{\prime}},d_{\alpha^{\prime}})<(\epsilon_{\alpha},d_{\alpha})$ $\text{ iff }(\epsilon_{\alpha^{\prime}}<\epsilon_{\alpha}\text{ and }d_{\alpha^{\prime}}<_{\mathbb{D}}d_{\alpha})$
$\text{ iff }({\alpha^{\prime}}<\alpha\text{ and }d_{\alpha^{\prime}}<_{\mathbb{D}}d_{\alpha})$ $\epsilon_{i}\text{'s are increasing}$
$\text{ iff }d_{\alpha^{\prime}}<_{\mathbb{D}}d_{\alpha}$ by condition ($\triangle$4)
[…]
To see $\mathcal{B}$ is a base for $\dot{U}_{\beta}[G_{\beta}]$, it is enough to verify that the elements of the base described in $(1)$ were generated by $\mathcal{B}$. Let $(\gamma,d)\in(I\cap\beta)\times\mathbb{D}$ be arbitrary. Find $\alpha$ such that $\epsilon_{\alpha}>\gamma$ (which exists since $I^{*}$ is unbounded) and let $\alpha^{\prime}$ be such that $d=d_{\alpha^{\prime}}$. Let $\alpha^{*}>\alpha,\alpha^{\prime}$ such that $d<_{\mathbb{D}}d_{\alpha^{*}}$. Hence, $B_{(\epsilon_{\alpha^{*}},d_{\alpha^{*}})}$ is contained in $B_{\gamma,d}$. ∎
Remark 4.5.
Note that if we replace the assumption in Lemma 4.4 with $I\subseteq S^{*}_{\dot{U}}$ and ($\dagger$3.5), then Lemma 4.4 holds for $\beta\in S^{*}_{\dot{U}}\cap\lim(I)$. Indeed, the argument for $(2)$ is identical. To see $(1)$, first argue that $\{B_{(\gamma,d)}\mid\gamma\in I\cap\beta,d\in\mathbb{D}\}\subseteq\dot{U}_{\beta}[G_{\beta}]$: using ($\dagger$3.5), if $\gamma\in I\cap\beta$, then for all $d\in\mathbb{D}$, $B_{(\gamma,d)}\in\dot{F}_{\beta,d^{\prime}}[G_{<(\beta,d^{\prime})}]$ for any $d^{\prime}>d$. Since $\beta\in S^{*}_{\dot{U}}$, Claim 4.3 ensures that $\dot{F}_{\beta,d^{\prime}}[G_{<(\beta,d^{\prime})}]\subseteq\dot{U}_{\beta}[G_{\beta}]$. Conversely, if $x\in\dot{U}_{\beta}[G_{\beta}]$ we use Claim 4.3 again to see that $x\in\dot{F}_{(\alpha,d)}[G_{<(\alpha,d)}]$ for some $\alpha\in I\cap\beta$ and $d\in\mathbb{D}$. Thus $B_{(\alpha,d)}\subseteq^{*}x$.
The point of Lemma 4.4 is that in order to finish the argument, it suffices to find an ultrafilter $U$ in the generic extension such that there are generic Mathias sets $B_{a}$ belonging to $U$ for sufficiently many $a\in\mathbb{D}^{+}$ such that condition ($\dagger$3.5) is satisfied.
The ultimate goal will be to have enough copies of $\mathbb{D}$ (i.e. fibers of $\mathbb{D}^{+}$) so that cofinally we can get simultaneously a copy of $\mathbb{D}$ which is cofinal. In order to obtain it, we will use a different argument for the countable case and the supercompact case.
4.1. At a measurable cardinal
Here we will prove the main consistency result for measurable cardinals.
@@ 4.1. At a measurable cardinal
Let $\kappa$ be supercompact cardinal. Then for every well-founded, $\kappa^{+}$-directed set $\mathbb{D}$ with a possible order-type $\mu$, there is a
$\text{(}\kappa\text{-cc) }\ast\text{ (}\mu^{+}\text{-directed closed) }\ast\text{ (}\kappa\text{-directed closed+}\kappa^{+}\text{-cc)}$
forcing extension in which there is a simple $P_{\mathbb{D}}$ ultrafilter over $\kappa$.
@@
To illustrate the main idea in a simple setting, suppose that $\kappa$ is a Laver indestructible supercompact cardinal and consider a linear iteration of ultrafilter Mathias forcing at $\kappa$ with $\kappa-$supports, where at stage $\alpha$ we use a measure $U_{\alpha}\in V[G_{\alpha}]$ to define the next step in the iteration: there is no problem in choosing such $U_{\alpha}$ because the forcing so far is $\kappa$-directed closed. In the final model $V[G]$ there is of course a measure $U$, but a priori we have no reason to believe that the set $B_{\alpha}$ added at stage $\alpha$ is in $U$, or even that $U_{\alpha}\subseteq U$.